OpenStudy (anonymous):

If f(x)=sin^2(3-x), then what is the exact value of f'(3-2pi/3)?

OpenStudy (anonymous):

\[ f'(x)=-2 \sin (3-x) \cos (3-x) \] Can you finish it now?

OpenStudy (anonymous):

\[ f'\left(3-\frac{2 \pi }{3}\right)=\frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

Ask me if you still do not understand it?